### 26 Jan 2024 v1

@@ 6 Implications of the environment on the evolution of the bar fraction
The bar fraction, defined as the ratio of disc galaxies having a bar to the total number of disc galaxies at a specific redshift and its evolution, has the potential to provide insights into the formation and evolution of bars and the assembly of the host galaxies. Several works have looked into the bar fraction, either in observations (e.g., Sheth et al. 2008; Gavazzi et al. 2015; Melvin et al. 2014; Cervantes-Sodi 2017; Erwin 2019) and simulations (e.g. Peschken & Lokas 2019, Fragkoudi et al. 2020, Reddish et al. 2022). In Rosas-Guevara et al. (2022), we look into the evolution of the bar fraction in the TNG50. There, we found a gradual evolution in the bar fraction with redshift, increasing from $0.28$ at $z=4$ and reaching the highest value of $0.48$ at $z=1$. Subsequently, it exhibits a smooth decline to $0.30$ at, $z=0$.
To understand the evolution of the bar fraction, we now look into the evolution of disc galaxies between $z=1$ to $z=0$. We characterise this evolution in Fig. 18 where we found:
@@
  • The morphology of the remainder changes to that of lenticulars, with $26.9$ per cent of them holding a bar and $1.1$ per cent not doing so.
  • $33.3$ per cent of the barred disc galaxies eventually transform into spheroids.
  • $38.7$ per cent of unbarred galaxies developed a bar at a later time.
  • $22.5$ per cent of unbarred galaxies never developed a bar.
  • $19.8$ per cent and $18.9$ per cent, respectively, transform into lenticulars and spheroids, respectively.
  • $4.5$ per cent of lenticulars fail to develop; and the remaining $14.4$ per cent maintains a bar at $z=0$.
Figure 18: Pie chart summarising the fate of barred and unbarred galaxies from $z=1$ to $z=0$. More than 50 per cent of disc galaxies suffer a morphological transformation whereas stable bars once formed, just a small percentage lose their bar.
Taking into account our findings, we can define the bar fraction, $f_{{\rm bar},zi}$ at a given redshift $z={i}$ as
@@
where, $n_{{\rm bar},zi}$ and $n_{{\rm disc},zi}$ are the number of disc galaxies with a bar and the number of discs at any given redshift, respectively. In the particular case of the bar fraction at $z=0$, it can be computed as follows:
@@
$=$ $f_{{\rm bar},z1}+\sum_{z=1}^{z=0}\frac{\Delta f_{{\rm bar},z1}}{\Delta t}\Delta t.$ (5)
The second term, which is the change of the bar fraction between $z=1$ and $z=0$, is contingent upon the formation and dissolution of bars during this period of time. This term can be subdivided into six distinct components, including the morphological transformation of galaxies, as outlined below:
$\sum_{z=1}^{z=0}\frac{\Delta f_{{\rm bar},z1}}{\Delta t}\Delta t$ $=$ $f_{{\rm barform},z<1}+f_{{\rm barform,discform}z<1}$
@@
The initial two terms exhibit positive values and are indicative of the emergence of bars after $z=1$. Conversely, the subsequent terms have negative values, representing the dissolution of a bar or the destruction of discs.
@@
The first term corresponds to the fraction of bars found on one of our disc descendant subsamples. More precisely, the subsample corresponds to the descendant galaxies that have a disc with a dominant cold component at $z=1$, but the bar forms after $z=1$. The given fraction can be represented as $f_{{\rm barform},z<1}=(n_{{\rm bar,form}}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0})$ where $n_{{\rm disc},z0}=349$ represents the number of discs at $z=0$ in the catalogue Rosas-Guevara et al. (2022); $n_{{\rm disc},z1}=204$ represents the number of galaxies with a cold component, which corresponds to the number of galaxies in our parent sample (see Table 1) and $n_{{\rm bar,form}}=43$ the number of bars that formed after $z<1$ of this subsample of descendant discs (see Table 1). The value of $f_{{\rm barform},z<1}$ can be calculated as $43/204\times 204/349=0.1$. It is worth noting that a portion of this value, approximately $30$ per cent, is attributed to environmental factors, specifically major/minor mergers. This observation is supported by the findings presented in subsection 5.1 and illustrated in Fig. 14 and Fig. 15. These figures indicate that the contribution of environmental factors to the formation of the bar fraction ($f_{{\rm barform},z<1}$) is limited to a maximum of $6$ per cent.
@@
The second term, $f_{{\rm barform,discform}z<1}$, corresponds to the bars formed in discs that have finished assembling the large cold disc component during a redshift lower than $1$. The expression may be represented as $f_{{\rm barform,discform}z<1}=(n_{{\rm bar,form}<z1}/n_{{\rm disc},z<1})(n_{{\rm disc},z<1}/n_{{\rm disc},z0})$ with $n_{{\rm disc},z<1}=349-204=145$. This is the number of discs that the cold disc component formed after $z=1$. The number of bars formed in these discs is represented as $n_{{\rm barform}<z1}=105-74$. The first number represents the number of bars observed in disc galaxies at $z=0$ and comes from the catalogue of Rosas-Guevara et al. (2022) (see their Table 2) and the second number is the number of bars of our disc descendants subsamples and is given in Table 1. This gives us the value of $f_{{\rm barform,discform}z<1}=0.1$. Assuming that $30$ per cent of the bars formed are a consequence of mergers (the same fraction identified in bars formed in the descendants of discs at $z=1$), it can be inferred that approximately $6$ per cent of the bars in the discs may be attributed to a major/minor merger event. This implies that a maximum of $12$ per cent of the bars at $z=0$, that formed after $z=1$, can be attributed to environmental influences. It should be noted that the aforementioned value of $12$ per cent serves as an upper bound. This is due to the assumption that the rate of mergers for the discs produced after $z=1$ remains constant. However, it is important to acknowledge that this assumption is not entirely accurate, since it is well-established that the merger rate for a given mass falls over time (see Fig 4 in Rodriguez-Gomez et al. 2015).
@@
The final four terms in Eq. LABEL:eq:changebarfraction are associated with the process of bar dissolution and exhibit negative values. The third term, $f_{{\rm barlost},z<1}$, represents the fraction of bars that have dissolved in barred disc galaxies at $z=1$. This corresponds to our subsample of disc galaxies that were barred at $z=1$ and afterwards experienced the dissolution of said bar. The calculation involves the product of two ratios $f_{{\rm barlost},z<1}=(n_{{\rm bar,lost}}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0})=5/204\times 204/349=0.01$. The first ratio, $(n_{\rm bar,lost}/n_{{\rm disc},z1})$, represents the number of galaxies with dissolved bars divided by the number of disc galaxies at $z=1$. The second ratio is the number of discs at $z=1$ over the number of discs at $z=0$. As we have seen in section 4.1 and Figs. 9 & 10, these galaxies suffer tidal influence from a massive companion in the temporal proximity to the dissolution of the bar. Additionally, it is noted that these galaxies tend to inhabit environments characterised by higher density. Considering the available evidence, it may be inferred that the aforementioned $1$ per cent can be attributed to environmental influences.
@@
The fourth term of Eq. LABEL:eq:changebarfraction, $f_{{\rm bar,desdisc},z<1}$, represents the fraction of bars in the descendants of disc galaxies at $z=1$ that experience damage or destruction of their cold disc component, resulting in a morphological transition. We can express $f_{{\rm bar,desdisc},z<1}=(n_{{\rm bars},z1}/n_{{\rm nondisc},z0})(n_{{\rm nondisc},z0}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0})$. The fraction ca be calculated by finding the number of barred galaxies in our sample that transform into lenticular and spheroidal, which corresponds to $26$ and $31$ respectively (see Table 1) and the number of the total galaxies that transform into lenticulars and spheroidals ($47$ and $53$). The combined values in $f_{{\rm bar,desdiscs},z<1}$ are determined to be $0.16$.
@@
The fifth term, denoted as $f_{{\rm barlost,disc,form},z<1}$, is aligned with the third term and represents the fraction of dissolved bars in discs where the formation of their cold disc component occurs after $z=1$. The given expression may be represented as $f_{{\rm barlost,disc,form},z<1}=(n_{{\rm bar,lost,discform}z<1}/n_{{\rm disc,form},z<1})$ $(n_{{\rm disc,form},z<1}/n_{{\rm disc},z0})$. The analysis conducted here does not incorporate the dissolution of the bars in the discs that formed a dominant cold component beyond $z=1$. However, it is reasonable to assume that similar processes, such as environmental factors, may contribute to this dissolution with comparable efficiency. Consequently, the ratio $(n_{{\rm bar,lost,disc,form}z<1}/n_{{\rm disc,form},z<1})$ is equal to $0.01$ and $(n_{{\rm disc,form},z<1}/n_{{\rm disc},z0})=145/349$. By substituting the appropriate values, we determined that $f_{{\rm barlost,disc,form},z<1}=0.004$.
@@
Lastly, the final term, $f_{{\rm disc,form,des},z<1}$, represents the fraction of discs that underwent the formation of their cold component after $z=1$, forming a bar and then experiencing the destruction of the disc. We assume that the efficiency of these transient discs is comparable to the one in our sample at $z=1$, and the fraction of disc galaxies undergoing morphological transformations is similar too. The calculated value for $f_{{\rm bar,disc,form,des},z<1}$ will be aligned with the fourth term, which is $0.16$.
Gathering our results, we have that Eq. LABEL:eq:changebarfraction is estimated as
@@
$=$ $-0.13$
The bar fraction of our parent sample at $z=1$ is $0.45\pm 0.03$ and based on this analysis and Eq. 5, the bar fraction that we obtained at $z=0$ is $0.31$, whereas the bar fraction at $z=0$ is $0.30\pm 0.02$, in good agreement, and the possible discrepancy that we have could be about the assumptions in the discs that formed later than $z=1$. Note also that this gave us information about the galaxies that assembled their massive cold disc component, after $z=1$ since the number of discs with a dominant cold component increases with decreasing redshift (see Table 1 in Rosas-Guevara et al. 2022 and Pillepich et al. 2019; van der Wel et al. 2014). This gives us a hint that maybe the net efficiency of forming bars is lower in discs that assembled later in time since there is a slight decrease in the bar fraction.
@@ 6.1 Outlook
The simulation TNG50 has allowed us to estimate the importance of the environment in the evolution of the bar fraction. It is important, however, to acknowledge the constraints associated with this estimation. For instance, we focus on the bar fraction from $z=1$ to $z=0$ as well as for massive disc galaxies ($M_{*}>10^{10}\rm M_{\sun}$). Nevertheless, we find bars in massive galaxies since $z=4$ in the simulation (Rosas-Guevara et al. 2022), which aligns with recent observations from the JWST at such high redshifts (Guo et al. 2023; Tsukui et al. 2023; Le Conte et al. 2023). The conditions at which galaxies inhabit at such high redshifts are notably distinct from lower redhsifts, where it is anticipated that the frequency of encounters and mergers is higher. In light of this, it is imperative to investigate the influence of the surrounding environment on the bar fraction at lower stellar masses and higher redshifts as has been suggested in Costantin et al. (2023) which found a bar in a low-mass disc galaxy ($3.9\times 10^{9}\rm M_{\sun}$ at $z=3$). Indeed, Zana et al. (2022) find a small bar fraction(less than 0.1, see their Fig. 17) in TNG50 disc galaxies with stellar mass $10^{9}$-$5\times 10^{9}\rm M_{\sun}$. However, the authors do not conduct an exhaustive investigation to determine whether these bars represent transient structures or exhibit temporal stability. Alternatively, further investigations into low-mass disc galaxies, as reported in Costantin et al. (2023) could give us new constraints in the bar fraction in disc galaxies in the low-mass range.
@@
In addition, the present analysis does not consider the potential weakening or disappearance of bars due to vertical instabilities in discs formed after $z=1$. However, earlier research utilising N-body simulations has demonstrated the ability of bars to reform under such circumstances (Bournaud, Combes, & Semelin 2005; Berentzen et al. 2004).
@@
An additional factor that we do not see in discs formed at $z=1$ but could affect a disc formed after $z=1$ is the disolution of bars due to the tidal interactions resulting from minor mergers. An illustrative instance can be observed in the study conducted by Ghosh et al. (2021), wherein N-body simulations were employed to investigate the impact of tidal interactions resulting from minor mergers on the central stellar bar in the remnant of the merger. Following each pericentre transit of the satellite, the central bar undergoes temporary phases of bar amplification. However, the primary occurrence of bar weakening occurs only after the merger event. Nevertheless, it is plausible that a galaxy could undergo the process of accreting cold gas either during mergers or at a subsequent stage, hence potentially revitalising the bar structure. In the conducted study, it was shown that minor mergers did not provide any discernible influence. However, it is worth noting that this phenomenon may hold significance in scenarios characterised by lower redshift or high-density conditions, where galaxies are subject to more exposed tidal interactions.
@@
Tidal interactions may play a significant role in the creation and dissolution of bars in galaxies with lower stellar masses and in highly dense environments. For example, the studies conducted by Łokas et al. (2014) and Łokas et al. (2016) examine the process of bar evolution in dwarf galaxies and Milky Way galaxies resulting from tidal interactions within a cluster-like setting. The authors discovered that the characteristics of the bars undergo temporal variations and are subject to the influence of the magnitude of the tidal force encountered during the evolutionary process. The formation of bars in galaxies is observed to occur at earlier stages and exhibits greater strength and length in galaxies that experience a higher degree of tidal force exerted by the surrounding cluster. This particular form of interaction has not been extensively examined; however, it is our intention to explore it in a forthcoming research project.
@@
AGN feedback has been identified as an additional mechanism linked to the formation and dissolution of a bar. In particular, Bonoli et al. (2016) find that a simulated galaxy forms a strong bar below $z\sim 1$, and the authors point out that the disc in the simulation is more prone to instabilities compared to the original Eris, possibly because of early AGN feedback affecting the central part of the galaxy. Zana et al. (2019), studying an enhanced suite of Eris, highlight the effects of the feedback processes on the formation time and final properties of the bar. In a study conducted by Łokas (2022), it was observed that a significant merger event resulting in the coalescence of two supermassive black holes led to an enhancement of the AGN feedback within the TNG 100 simulation. This, in turn, caused the expulsion of gas from the central region of the galaxy, subsequently suppressing star formation and facilitating the formation of a bar structure. However, it is worth noting that AGN feedback may potentially exert a counteractive influence by weakening the bar and potentially resulting in their eventual removal (Irodotou et al. 2022). This process is a potential channel to understand the formation and dissolution of a bar. We have identified a specific instance in which, after a significant interaction with a companion, a notable episode of strong AGN feedback, following the disappearance of the bar. Nevertheless, it remains uncertain whether the tidal interaction is responsible for the disappearance of the bar and the subsequent activation of the AGN, or if the AGN feedback itself plays a role in the dissolution of the bar.
@@
One crucial consideration to be mindful of is the constrained volume of TNG50, which presents limitations in terms of its ability to comprehensively investigate a wide range of diverse environments, particularly for disc galaxies in a wider range of stellar masses. This is another aspect to see detail in future with the next generation of cosmological hydrodynamical simulations.

### 07 Feb 2024 v2

@@ 6 Implications of the environment on the evolution of the bar fraction
The bar fraction —defined as the ratio of the number of disc galaxies with a bar to the total number of disc galaxies at a specific redshift— and its evolution has the potential to provide insights into the formation and evolution of bars and the assembly of the host galaxies. Several works have looked into the bar fraction, either in observations (e.g. Sheth et al. 2008; Gavazzi et al. 2015; Melvin et al. 2014; Cervantes-Sodi 2017; Erwin 2019) or simulations (e.g. Peschken & Lokas 2019, Fragkoudi et al. 2020, Reddish et al. 2022). In Rosas-Guevara et al. (2022), we look into the evolution of the bar fraction in the TNG50. There, we find a gradual evolution in the bar fraction with redshift, increasing from $0.28$ at $z=4$ and reaching the highest value of $0.48$ at $z=1$. Subsequently, the bar fraction exhibits a smooth decline to $0.30$ at $z=0$.
To understand this evolution, we now look into the evolution of the bar in disc galaxies between $z=1$ and $z=0$. We characterise this evolution in Fig. 18, where we find:
@@
  • The morphology of the remaining galaxies changes to that of lenticular galaxies, with $26.9$ per cent of them holding a bar and $1.1$ per cent not holding one.
  • $33.3$ per cent of the barred disc galaxies eventually transform into spheroid galaxies.
  • $38.7$ per cent of unbarred galaxies develop a bar at a later time.
  • $22.5$ per cent of unbarred galaxies never develop a bar.
  • $19.8$ per cent and $18.9$ per cent transform into lenticular and spheroid galaxies, respectively.
  • $4.5$ per cent of lenticular galaxies fail to develop; and the remaining $14.4$ per cent maintain a bar at $z=0$.
Figure 18: Pie chart summarising the fate of barred and unbarred galaxies from $z=1$ to $z=0$. More than 50 per cent of disc galaxies suffer a morphological transformation, whereas just a small percentage of galaxies with a stable bar, lose their bar.
Taking into account our findings, we can define the bar fraction, $f_{{\rm bar},zi}$ , at a given redshift $z={i}$ as
@@
where $n_{{\rm bar},zi}$ and $n_{{\rm disc},zi}$ are the number of disc galaxies with a bar and the number of discs at any given redshift, respectively. In the particular case of the bar fraction at $z=0$, it can be computed as follows:
@@
$=$ $f_{{\rm bar},z1}+\sum_{z=1}^{z=0}\frac{\Delta f_{{\rm bar},z1}}{\Delta t}\Delta t$ (5)
The second term, which is the change of the bar fraction between $z=1$ and $z=0$, is contingent on the formation and dissolution of bars during this period of time. This term can be subdivided into six distinct components, including the morphological transformation of galaxies, as outlined below:
$\sum_{z=1}^{z=0}\frac{\Delta f_{{\rm bar},z1}}{\Delta t}\Delta t$ $=$ $f_{{\rm barform},z<1}+f_{{\rm barform,discform}z<1}$ (6)
@@
The initial two terms exhibit positive values and are indicative of the emergence of bars after $z=1$. Conversely, the subsequent terms have negative values, which indicate the dissolution of a bar or the destruction of a disc.
@@
The first term of Eq. 6 corresponds to the fraction of bars found on one of our disc descendant subsamples. More precisely, the subsample corresponds to the descendant galaxies that have a disc with a dominant cold component at $z=1$, but with a bar forming after $z=1$. The given fraction can be represented as $f_{{\rm barform},z<1}=(n_{{\rm bar,form}}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0}),$ where $n_{{\rm disc},z0}=349$ represents the number of discs at $z=0$ in the catalogue of Rosas-Guevara et al. (2022); $n_{{\rm disc},z1}=204$ represents the number of galaxies with a cold component, which corresponds to the number of galaxies in our parent sample (see Table 1) and $n_{{\rm bar,form}}=43$ the number of bars that formed after $z<1$ of this subsample of descendant discs (see Table 1). The value of $f_{{\rm barform},z<1}$ can be calculated as $43/204\times 204/349=0.1$. It is worth noting that a fraction of this value, namely approximately $30$ per cent, is attributed to environmental factors, specifically major or minor mergers. This observation is supported by the findings presented in subsection 5.1 and illustrated in Fig. 14 and Fig. 15. These figures indicate that the contribution of environmental factors (mergers) to the formation of new bars in the bar fraction ($f_{{\rm barform},z<1}$) is limited to a maximum of $6$ per cent.
@@
The second term, $f_{{\rm barform,discform}z<1}$, corresponds to the bars formed in discs that have finished assembling the large cold-disc component at a redshift of lower than $1$. The expression may be represented as $f_{{\rm barform,discform}z<1}=(n_{{\rm bar,form}<z1}/n_{{\rm disc},z<1})(n_{{\rm disc},z<1}/n_{{\rm disc},z0}),$ where $n_{{\rm disc},z<1}=349-204=145$. This is the number of discs that the cold-disc component formed after $z=1$. The number of bars formed in these discs is represented as $n_{{\rm barform}<z1}=105-74$. The first number represents the number of bars observed in disc galaxies at $z=0$ and comes from the catalogue of Rosas-Guevara et al. 2022, see their Table 2 and the second number is the number of bars of our disc descendant subsamples and is given in Table 1. This gives us the value of $f_{{\rm barform,discform}z<1}=0.1$. Assuming that $30$ per cent of the bars formed are a consequence of mergers (the same fraction identified in bars formed in the descendants of discs at $z=1$), it can be inferred that approximately $6$ per cent of the bars in the discs may be attributed to a major or minor merger event. This implies that a maximum of $12$ per cent of the bars at $z=0$ that formed after $z=1$ can be attributed to environmental influences. It should be noted that the aforementioned value of $12$ per cent serves as an upper bound. This is due to the assumption that the rate of mergers for the discs produced after $z=1$ remains constant. However, it is important to acknowledge that this assumption is not entirely accurate, as it is well established that the merger rate for a given mass falls over time (see Fig 4 in Rodriguez-Gomez et al. 2015).
@@
The final four terms in Eq. 6 are associated with the process of bar dissolution and exhibit negative values. The third term, $f_{{\rm barlost},z<1}$, represents the fraction of bars that have dissolved in barred disc galaxies at $z=1$. This corresponds to our subsample of disc galaxies that were barred at $z=1$ and afterwards experienced the dissolution of that bar. The calculation involves the product of two ratios $f_{{\rm barlost},z<1}=(n_{{\rm bar,lost}}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0})=5/204\times 204/349=0.01$. The first ratio, $(n_{\rm bar,lost}/n_{{\rm disc},z1})$, represents the number of galaxies with dissolved bars divided by the number of disc galaxies at $z=1$. The second ratio is the number of discs at $z=1$ over the number of discs at $z=0$. As we have seen in section 4.2 and Figs. 9 and 10, these galaxies suffer a tidal influence from a massive companion in temporal proximity to the dissolution of the bar. Additionally, it is noted that these galaxies tend to inhabit environments characterised by higher density. Considering the available evidence, it may be inferred that the aforementioned $1$ per cent can be attributed to environmental influences.
@@
The fourth term of Eq. 6, $f_{{\rm bar,desdisc},z<1}$, represents the fraction of bars in the descendants of disc galaxies at $z=1$ that experience damage or destruction of their cold-disc component, resulting in a morphological transition. We can express $f_{{\rm bar,desdisc},z<1}=(n_{{\rm bars},z1}/n_{{\rm nondisc},z0})(n_{{\rm nondisc},z0}/n_{{\rm disc},z1})(n_{{\rm disc},z1}/n_{{\rm disc},z0})$. This fraction can be calculated by finding the number of barred galaxies in our sample that transform into lenticular and spheroidal galaxies, which corresponds to $26$ and $31,$ respectively (see Table 1) and the number of the total galaxies that transform into lenticulars and spheroidals ($47$ and $53$). The combined values in $f_{{\rm bar,desdiscs},z<1}$ are determined to be $0.16$.
@@
The fifth term, denoted $f_{{\rm barlost,disc,form},z<1}$, is aligned with the third term and represents the fraction of dissolved bars in discs where the formation of their cold-disc component occurs after $z=1$. The given expression may be represented as $f_{{\rm barlost,disc,form},z<1}=(n_{{\rm bar,lost,discform}z<1}/n_{{\rm disc,form},z<1})$ $(n_{{\rm disc,form},z<1}/n_{{\rm disc},z0})$. The analysis conducted here does not incorporate the dissolution of the bars in the discs that formed a dominant cold component beyond $z=1$. However, it is reasonable to assume that similar processes, such as environmental factors, may contribute to this dissolution with comparable efficiency. Consequently, the ratio $(n_{{\rm bar,lost,disc,form}z<1}/n_{{\rm disc,form},z<1})$ is equal to $0.01$ and $(n_{{\rm disc,form},z<1}/n_{{\rm disc},z0})=145/349$. By substituting the appropriate values, we determine that $f_{{\rm barlost,disc,form},z<1}=0.004$.
@@
Lastly, the final term, $f_{{\rm disc,form,des},z<1}$, represents the fraction of discs that underwent the formation of their cold component after $z=1$, forming a bar and then experiencing the destruction of the disc. We assume that this fraction is comparable to that of the descendant galaxies whose cold-disc component undergoes damage or complete destruction, leading to a morphological transition. In such cases, the calculated value for $f_{{\rm bar,disc,form,des},z<1}$ will be aligned with the fourth term of Eq. 6, which is $0.16$.
Gathering our results, we estimate the product of Eq. 6 to be as follows:
@@
$=$ $-0.13.$
The bar fraction of our parent sample at $z=1$ is $0.45\pm 0.03$ and based on this analysis and Eq. 5, the bar fraction that we obtain at $z=0$ is $0.31$, whereas the bar fraction at $z=0$ is $0.30\pm 0.02$, which is in good agreement, and the possible discrepancy that we have could be linked to the assumptions for the discs that formed later than $z=1$. We also note that this gave us information about the galaxies that assembled their massive cold-disc component after $z=1,$ because the number of discs with a dominant cold component increases with decreasing redshift (see Table 1 in Rosas-Guevara et al. 2022 and Pillepich et al. 2019; van der Wel et al. 2014). This suggests that the net efficiency of forming bars could be lower in discs that assembled later in time, because there is a slight decrease in the bar fraction.
@@ 6.1 Outlook
The simulation TNG50 allows us to estimate the importance of the environment in the evolution of the bar fraction. However, we must acknowledge the constraints associated with this estimation. For instance, we focus on the bar fraction from $z=1$ to $z=0$ as well as for massive disc galaxies ($M_{*}>10^{10}\rm M_{\sun}$). Nevertheless, we find bars in massive galaxies since $z=4$ in the simulation (Rosas-Guevara et al. 2022), which aligns with recent observations from the JWST at such high redshifts (Guo et al. 2023; Tsukui et al. 2023; Le Conte et al. 2023). The conditions experienced by galaxies at such high redshifts are notably distinct from those at lower redshifts, where it is anticipated that the frequency of encounters and mergers is higher. In light of this, it is imperative to investigate the influence of the surrounding environment on the bar fraction at lower stellar masses and higher redshifts, as has been suggested by Costantin et al. (2023), who, for example, found a bar in a low-mass disc galaxy ($3.9\times 10^{9}\rm M_{\sun}$ at $z=3$). Indeed, Zana et al. (2022) find a small bar fraction (less than 0.1, see their Fig. 17) in TNG50 disc galaxies with stellar masses of $10^{9}$$5\times 10^{9}\rm M_{\sun}$. However, the authors do not conduct an exhaustive investigation to determine whether these bars represent transient structures or exhibit temporal stability. Alternatively, further investigations into low-mass disc galaxies —as reported in Costantin et al. (2023)— could give us new constraints of the bar fraction in disc galaxies in the low-mass range.
@@
In addition, the present analysis does not consider the potential weakening or disappearance of bars due to vertical instabilities in discs formed after $z=1$. However, earlier research using N-body simulations demonstrated the ability of bars to reform under such circumstances (Bournaud, Combes, & Semelin 2005; Berentzen et al. 2004).
@@
An additional factor that we do not see in discs formed at $z=1$ but that could affect a disc formed after $z=1$ is the dissolution of bars due to the tidal interactions resulting from minor mergers. An illustrative instance can be observed in the study conducted by Ghosh et al. (2021), wherein N-body simulations were employed to investigate the impact of tidal interactions resulting from minor mergers on the central stellar bar in the remnant of the merger. Following each pericentre transit of the satellite, the central bar undergoes temporary phases of bar amplification. However, the primary occurrence of bar weakening occurs only after the merger event. Nevertheless, it is plausible that a galaxy could undergo the process of accreting cold gas either during mergers or at a subsequent stage, thus potentially revitalising the bar structure. Ghosh et al. (2021) showed that minor mergers do not provide any discernible influence. However, it is worth noting that this phenomenon may hold significance in scenarios characterised by lower redshift or high-density conditions, where galaxies are subject to more exposed tidal interactions.
@@
Tidal interactions may play a significant role in the creation and dissolution of bars in galaxies with lower stellar masses and in highly dense environments. For example, the studies conducted by Łokas et al. (2014) and Łokas et al. (2016) examine the process of bar evolution in dwarf galaxies and Milky Way galaxies resulting from tidal interactions within a cluster-like setting. The authors discovered that the characteristics of the bars exhibit temporal variations and are subject to the influence of the magnitude of the tidal force encountered during the evolutionary process. The formation of bars in galaxies is observed to occur at earlier stages and the bars themselves are longer and stronger in galaxies that experience a higher degree of tidal force exerted by the surrounding cluster. This particular form of interaction has not been extensively examined; however, it is our intention to explore this topic in a forthcoming research project.
@@
AGN feedback has been identified as an additional mechanism linked to the formation and dissolution of a bar. In particular, Bonoli et al. (2016) find that a simulated galaxy forms a strong bar below $z\sim 1$, and the authors point out that the disc in the simulation is more prone to instabilities compared to the original Eris, possibly because of early AGN feedback affecting the central part of the galaxy. Zana et al. (2019), studying an enhanced suite of Eris, highlight the effects of the feedback processes on the formation time and final properties of the bar. In a study conducted by Łokas (2022), it was observed that a significant merger event resulting in the coalescence of two supermassive black holes led to an enhancement of the AGN feedback within the TNG 100 simulation. This, in turn, caused the expulsion of gas from the central region of the galaxy, subsequently suppressing star formation and facilitating the formation of a bar structure. However, it is worth noting that AGN feedback may potentially exert a counteractive influence by weakening the bar and potentially resulting in their eventual removal (Irodotou et al. 2022). This process is a potential channel to understand the formation and dissolution of a bar. We identified a specific instance in which a bar disappeared following a significant interaction with a companion, with a subsequent, notable episode of strong AGN feedback. Nevertheless, it remains uncertain as to whether the tidal interaction is responsible for the disappearance of the bar and the subsequent activation of the AGN, or if the AGN feedback itself plays a role in the dissolution of the bar.
@@
One crucial consideration to be mindful of is the constrained volume of TNG50, which presents limitations in terms of its ability to allow comprehensive investigations of a wide range of diverse environments, particularly for disc galaxies in a wider range of stellar masses. This is another aspect that may see considerable development in the future with the next generation of cosmological hydrodynamical simulations.